(2+3i)^2 could you please explain step by step
please help soon!

Answer by ankor@dixie-net.com(22740)

(2+3i)^2 could you please explain step by step
:
Assume you want to FOIL (2+3i)(2+3i)
Step by step
F: mult first two term 2 * 2 = 4
0: mult outer terms: 2 * 3i = 6i
I: mult inner terms: 3i * 2 = 6i
L: mult last terms: 3i * 3i = 9i^2
:
So we have:
4 + 6i + 6i + 9i^2
:
Add like terms
4 + 12i + 9i^2
;
Remember that i^2 = -1 so you have:
4 + 12i + 9(-1)
4 + 12i - 9
:
4 - 9 + 12i
:
which is:
-5 + 12i